10. Solve the equation below on the interval (0.2T). cos2x = V3
Answer #1

16) Cos 27 V2 we Can Split dhis into 2 Passible (ases V3/2 Cos (27) = Cos-! Cos 21-2X = 22 =-> 27-2x =76 periuds adding by 2x(2 +na, nez Since nez, Then -ng =na] - (1) interual 0,27) In In put 11 n+0 nlin => in Put n:1 137 Th 237 1)19 Put n= 2, ho Interval 6, 19) ared values the in Ther 137/12, 237/12) (in, Ali2i